A right circular cone is inscribed in a cube having an edge which measures 20 cm. Find the lateral area and volume of the cone


1
Expert's answer
2021-01-11T10:00:50-0500

Solution


For cone inscribed in a cube with edge a=20 cm its diameter d=a and height h=a.

So volume V = (1/3)*AB*h = (1/3)*(π*d2/4)*h = π*d2*h/12 = π*a3/12 = π*203/12 = π*2000/3 ≈ 2094.4 cm3  (AB is area of the cone base).

The slant height of a right circular cone "L=\sqrt{(d/2)^2+h^2}=a\sqrt{5}/2"

Lateral area "LSA=\pi d L/2= \pi a^2 \sqrt{5}/4 \approx 702.48 cm^2"

Answer

volume V = π*a3/12 ≈ 2094.4 cm3

lateral area "LSA= \pi a^2 \sqrt{5}/4 \approx 702.48 cm^2"